Doggie Pawz produces two models of dog beds: rectangle and circle. The following information about the costs to produce the company’s products has been provided.

Category Estimated Cost Cost Driver Rectangle Circle
Unit-Level $293, 000 Labor Hours 200 300
Batch-Level $45, 000 Inspections 19 15
Product-Level $390,000 Storage Space 2900 sq. ft 3700 sq.ft
Facility-Level $125,000 Machine Hours 4,000 7,000

a. If all costs are allocated based on machine hours, what is the rate per machine hour?
Answer #1

Answer a.

Total Estimated Cost = $293000+$45000+$390000+$125000 = $853000

Total Machine Hours = 4000+7000 = 11000 hours

Machine Hour Rate = $853000/11000 = $77.5455 or 77.55 (rounded off to two decimals)

Hit Thumbs up if satisfied

Thank you